# Adaptive Depth Networks with Skippable Sub-Paths

### v1

@@ Abstract
Systematic adaptation of network depths at runtime can be an effective way to control inference latency and meet the resource condition of various devices. However, previous depth adaptive networks do not provide general principles and a formal explanation on why and which layers can be skipped, and, hence, their approaches are hard to be generalized and require long and complex training steps. In this paper, we present an architectural pattern and training method for adaptive depth networks that can provide flexible accuracy-efficiency trade-offs in a single network. In our approach, every residual stage is divided into 2 consecutive sub-paths with different properties. While the first sub-path is mandatory for hierarchical feature learning, the other is optimized to incur minimal performance degradation even if it is skipped. Unlike previous adaptive networks, our approach does not iteratively self-distill a fixed set of sub-networks, resulting in significantly shorter training time. However, once deployed on devices, it can instantly construct sub-networks of varying depths to provide various accuracy-efficiency trade-offs in a single model. We provide a formal rationale for why the proposed architectural pattern and training method can reduce overall prediction errors while minimizing the impact of skipping selected sub-paths. We also demonstrate the generality and effectiveness of our approach with various residual networks, both from convolutional neural networks and vision transformers.
@@ 1 Introduction
There have been some attempts to provide predictable adaptability to neural networks by exploiting the redundancy in either network depths [24, 6], widths [59, 48], or both [47, 20]. However, one major difficulty with prior adaptive networks is that they are hard to train and require significantly longer training time than non-adaptive networks. For example, most adaptive networks select a fixed number of sub-networks of varying depths or width, and train them iteratively, mostly through self-distilling knowledge from the largest sub-network (also referred to as the super-net) [59, 48, 20, 43]. However, this iterative self-distillation takes long time and can generate conflicting training objectives for different parameter-sharing sub-networks, potentially resulting in worse performance [30, 10]. Further, unlike width-adaptation networks, no general principle has been proposed for how to select sub-networks for depth adaptation, since the effect of skipping individual layers has not been formally specified.
@@
Figure 1: (a) At training time, the last half layers of every residual stage is optimized to have minimal performance degradation even if they are skipped. No explicit self-distillation is required for sub-networks, except the smallest base-net. (b) At test time, $2^{N_{r}}$ sub-networks become available in a single network by systematically skipping last half layers of $N_{r}$ residual stages. (c) Sub-networks instantly selected from a single network outperform counterpart individually trained networks due to regularization effect.
@@
In this work, we introduce an architectural pattern and training method for adaptive depth networks that is generally applicable to residual networks, e.g., CNNs and transformers. In the proposed adaptive depth networks, every residual stage is divided into 2 consecutive sub-paths and they are trained to have different properties. While the first sub-paths are mandatory for hierarchical feature learning, the second sub-paths are optimized to incur minimal performance degradation even if they are skipped. More specifically, the second sub-path of every residual stage is optimized to preserve the feature distribution from its previous mandatory sub-path in order to minimize the performance degradation when it is skipped. During training, this property of the second sub-paths is enforced through skip-aware self-distillation, in which only one smallest sub-network, also referred as the base-net, is jointly trained in order to self-distill intermediate feature distributions at every residual stage, as shown in Figure-1-(a). This skip-aware self-distillation requires no explicit self-distillation or additional fine-tuning of individual sub-networks except the base-net, resulting in significantly shorter training time than previous adaptive networks. However, once trained, sub-networks with various depths can be selected instantly from a single network to meet the resource condition of devices, as shown in Figure 1-(b). Further, these sub-networks with various depths outperform individually trained non-adaptive networks due to the regularization effect, as shown in Figure 1-(c).
@@
In Section 3, we discuss the details of our architectural pattern and training algorithm, and show formally that the chosen sub-paths trained with our skip-aware self-distillation are optimized to reduce prediction errors while minimally changing the level of input features. In Section 4, we empirically demonstrate that our adaptive depth networks with skippable sub-paths outperform counterpart individual networks, both in CNNs and vision transformers, and achieve actual inference acceleration and energy-saving.
To authors’ best knowledge, this work is the first general approach to adaptive depth networks, providing a general principle for depth adaptation and a formal explanation on why layers can be skipped with minimal performance degradation.
Figure 2: Illustration of a residual stage with 2 sub-paths. While the first sub-path is mandatory for hierarchical feature learning, the second (yellow) sub-path can be skipped for efficiency. The layers in the skippable sub-path are trained to preserve the feature distribution from $\mathbf{h}^{s}_{base}$ to $\mathbf{h}^{s}_{super}$ using skip-aware self-distillation. Having similar feature distributions, either $\mathbf{h}^{s}_{base}$ or $\mathbf{h}^{s}_{super}$ can be provided as an input to the next residual stage. In the mandatory sub-path, another set of batch normalization operators, called skip-aware BNs, is exploited when the second sub-path is skipped. This architectural pattern repeats in every residual stage throughout the network.
@@ 2 Related Work
Adaptive Networks: In most adaptive networks, parameter-sharing sub-networks are selected by adjusting either widths, depths, or resolutions [24, 59, 48, 22, 20, 30, 61, 47, 3]. For example, slimmable neural networks adjust channel widths of CNN models on the fly for accuracy-efficiency trade-offs and they exploit switchable batch normalization to handle multiple sub-networks [59, 58, 48]. Transformer-based adaptive depth networks have been proposed for language models to dynamically skip some of the layers during inference [20, 6]. However, in these adaptive networks, every target sub-network with varying widths or depths need to be trained explicitly, incurring significant training overheads and potential conflicts between sub-networks.
Dynamic networks [15] are another class of adaptive networks that exploit additional control networks or decision gates for input-dependent adaption of CNN models [51, 29, 12, 31, 53] and transformers [35, 55, 8, 18]. In particular, most dynamic networks for depth-adaptation have some kinds of decision gates at every layers (or blocks) that determine if the layers can be skipped [9, 51, 45, 50]. These approaches are based on the thought that some layers can be skipped on ‘easy’ inputs. However, the learned policy for skipping layers is opaque to users and does not provide a formal description of when and which layers can be skipped for a given input. Therefore, the network depths cannot be adapted in a predictable manner to meet the resource condition of target devices.
Residual Blocks with Shortcuts: Since the introduction of ResNets [16], residual blocks with shortcuts have received extensive attention because of their ability to train very deep networks, and have been chosen by many CNNs [38, 41] and transformers [44, 5, 34]. In [46], Veit et al. argue that identity shortcuts make exponential paths and results in an ensemble of shallower sub-networks. This thought is supported by the fact that removing individual residual blocks at test time does not significantly affect performance, and it has been further exploited to train deep networks [23, 52]. Other works argue that identity shortcuts enable residual blocks to perform iterative feature refinement, where each block improves slightly but keeps the semantic of the representation of the previous layer [11, 27]. Our work build upon those views on residual blocks with shortcuts and further extend them for adaptive depth networks by introducing an architectural pattern and training method that exploits the properties of residual blocks more explicitly for selected sub-paths.
@@ 3 Adaptive Depth Networks
We first present an architectural pattern and training details to build adaptive depth networks. And, then, we discuss the theoretic rationale for how depth adaption can be achieved with minimal performance degradation.
@@ 3.1 Architectural Pattern for Depth Adaptation
In typical hierarchical residual networks such as ResNets [16] and Swin transformers [34], the $s$-th residual stage is consisted of $L$ identical residual blocks, which transform input features $\mathbf{h}^{s}_{1}$ additively to produce the output features $\mathbf{h}^{s}$, as follows:
$\underbrace{\underbrace{\mathbf{h}^{s}_{1}+\underbrace{F_{1}(\mathbf{h}^{s}_{1})+...+{F}^{s}_{L/2}(\mathbf{h}^{s}_{L/2})}_{\mathbf{F}^{s}_{base}}}_{=\mathbf{h}^{s}_{base}}+\underbrace{...+F_{L}(\mathbf{h}^{s}_{L})}_{\mathbf{F}^{s}_{skippable}}}_{=\mathbf{h}^{s}_{super}}=\mathbf{h}^{s}$ (1)
@@
While a block with a residual function $F_{\ell}$ ($\ell=1,...,L$) learns hierarchical features as traditional compositional networks [40], previous literature [27, 11] demonstrates that a residual function also tend to learn a function that refines already learned features at the same feature level. If a residual block mostly performs feature refinement while not changing the level of input features, the performance of the residual network is not significantly affected by dropping the block at test time [23, 52]. However, in typical residual networks, most residual blocks tend to refine features while learning new level features as well, and, hence, random dropping of residual blocks at test time degrades the performance significantly. Therefore, we hypothesize that if some designated residual blocks can be encouraged during training to focus more on feature refinement, then these blocks can be skipped to save computation at marginal loss of prediction accuracy at test time.
@@
To this end, we propose an architectural pattern for adaptive depth networks, in which every residual stage is divided into two consecutive sub-paths, or $\mathbf{F}^{s}_{base}$ and $\mathbf{F}^{s}_{skippable}$ as in Equation 1 and Figure 2. While $\mathbf{F}^{s}_{base}$ learns feature representation $\mathbf{h}^{s}_{base}$ (= $\mathbf{h}^{s}_{L/2+1}$) with no constraint, the second sub-path $\mathbf{F}^{s}_{skippable}$ is constrained to preserve the feature level of $\mathbf{h}^{s}_{base}$ and only refine it to produce $\mathbf{h}^{s}_{super}$. Since layers in $\mathbf{F}^{s}_{base}$ perform essential transformations for hierarchical feature learning, they cannot be bypassed during inference. However, since layers in $\mathbf{F}^{s}_{skippable}$ only refine $\mathbf{h}^{s}_{base}$, they can be skipped to save computation. If $\mathbf{F}^{s}_{skippable}$ is skipped, then intermediate features $\mathbf{h}^{s}_{base}$ becomes the input to the next residual stage. Therefore, the overall network depth can be adjusted by choosing whether to skip $\mathbf{F}^{s}_{skippable}(s=1,...,N_{s})$ or not.
In Section 4.1, we show that this architectural pattern for adaptive depth networks is generally applicable to a wide range of residual networks.
3.2 Skip-Aware Self-Distillation
@@
Preserving the feature level of $\mathbf{h}^{s}_{base}$ in $\mathbf{F}^{s}_{skippable}$ implies, more specifically, that two feature representations $\mathbf{h}^{s}_{base}$ and $\mathbf{h}^{s}_{super}$ have similar distributions over training input $\mathbf{X}$. If $\mathbf{h}^{s}_{super}$ and $\mathbf{h}^{s}_{base}$ have similar distributions, skipping the sub-path $\mathbf{F}^{s}_{skippable}$ during inference results in minimal internal covariate shifts [25] to the following network layers. Kullback-Leibler (KL) divergence measures how different two distributions are over the same random variable, and, hence, we use Equation 2 to measure the similarity of two distributions, $\mathbf{h}^{s}_{base}$ and $\mathbf{h}^{s}_{super}$, over input $\mathbf{X}$:
$\underset{x\in\mathbf{X}}{D_{KL}}(\mathbf{h}^{s}_{super}||\mathbf{h}^{s}_{base})$ (2)
Algorithm 1 shows our training method, called skip-aware self-distillation, in which Equation 2 is included in the loss function while the largest and the smallest sub-networks of $\mathbf{M}$, which are called super-net and the base-net, respectively, are jointly trained  [^1] Some frameworks, such as Pytorch’s distributed data parallel (DDP), do not support two consecutive forwards passes followed by a single backward pass: $super.forward\rightarrow base.fowrard\rightarrow loss.backward$. For such frameworks, we adapt the algorithm to $super.forward\rightarrow loss_{super}.backward\rightarrow base.forward\rightarrow loss_{sub\_path}.backward$. The latter method has similar results and is more memory-efficient. . In Algorithm 1, the forward function of the proposed adaptive depth networks $\mathbf{M}$ accepts an extra argument, ’skip’, that controls which residual stages skip their skippable sub-paths. For example, if $\mathbf{M}$ has 4 residual stages, its base-net is selected by passing ‘skip=[True, True, True, True]’.
Algorithm 1 Training algorithm for an adaptive depth network $\mathbf{M}$. Two parameter-sharing sub-networks, or super-net and base-net, are jointly trained using skip-aware self-distillation. $N_{r}$ denotes the number of residual stages.
@@
8:     $loss_{base}=\sum_{s=1}^{N_{r}}D_{KL}(\mathbf{h}^{s}_{super}\|\mathbf{h}^{s}_{base})\newline \hskip 50.00008pt+D_{KL}(\mathbf{\hat{y}}_{super}\|\mathbf{\hat{y}}_{base})$
9:     $loss=\alpha loss_{super}+(1-\alpha)loss_{base}$
10:     $loss.backward()$
@@
In steps 5 and 7, the forward passes of the super-net and the base-net are executed for the same input $\mathbf{x}$. Intermediate features $\mathbf{h_{super}}$ and $\mathbf{h_{base}}$ are also obtained during the forward passes of $\mathbf{M}$. In step 8, $D_{KL}(\mathbf{h}^{s}_{super}||\mathbf{h}^{s}_{base})$ in Equation 2 is included in the loss function $loss_{base}$. By minimizing $loss_{base}$, two feature representations $\mathbf{h}^{s}_{super}$ and $\mathbf{h}^{s}_{base}$, are explicitly enforced to have similar distributions for the same input $\mathbf{x}$. Further, this step gives the effect of transferring the knowledge from $\mathbf{h}^{s}_{super}$ to $\mathbf{h}^{s}_{base}$ at every residual stage. Therefore, $\mathbf{h}^{s}_{base}$ is expected to learn more compact representation from $\mathbf{h}^{s}_{super}$. The hyperparameter $\alpha$ controls the strength of this skip-aware self-distillation. In step 8, $D_{KL}(\mathbf{\hat{y}}_{super}\|\mathbf{\hat{y}}_{base})$ is included in the loss function for further distillation effect from the super-net to the sub-networks. Due to the architectural pattern of interleaving the mandatory and the skippable sub-paths, minimizing $D_{KL}(\mathbf{\hat{y}}_{super}\|\mathbf{\hat{y}}_{base})$ also minimizes $D_{KL}(\mathbf{h}_{super}\|\mathbf{h}_{base})$ implicitly. Experimental result shows that similar results can be achieved when only $D_{KL}(\mathbf{\hat{y}}_{super}\|\mathbf{\hat{y}}_{base})$ is used in $loss_{base}$ (Section 4.5). This implicit approach is useful when the extraction of intermediate features is tricky.
In Algorithm 1, only two sub-networks are involved for the training of $\mathbf{M}$, and, hence, the total training time is no greater than training two sub-networks individually. However, at test time, sub-networks with various depths can be selected on the fly by systematically skipping sub-paths in residual stages. For example, in a network with 4 residual stages, $16\ (=2^{4})$ parameter-sharing sub-networks can be selected by varying the skip argument. In contrast, prior adaptive networks supporting 16 parameter-sharing sub-networks need to perform 16 explicit self-distillation from the super-net to the sub-networks [48, 59, 20]. We demonstrate this result in Section 4.3.
3.3 Formal Analysis of Skippable Sub-Paths
$D_{KL}(\mathbf{h}^{s}_{super}||\mathbf{h}^{s}_{base})$ in Equation 2 can be trivially minimized if residual blocks in $\mathbf{F}^{s}_{skippable}$ learn identity functions, or $\mathbf{h}^{s}_{base}+\mathbf{F}^{s}_{skippable}(\mathbf{h}^{s}_{base})=\mathbf{h}^{s}_{base}$. However, since the super-net is jointly trained with the loss function $loss_{super}$, the residual functions in $\mathbf{F}^{s}_{skippable}$ cannot simply be an identity function. Then, what do the residual functions in $\mathbf{F}^{s}_{skippable}$ learn during training? This can be further investigated through Taylor expansion [27]. For our adaptive depth networks, a loss function $\mathcal{L}$ used for training the super-net can be approximated with Taylor expansion as follows:

### v2

@@ Abstract
Predictable adaptation of network depths can be an effective way to control inference latency and meet the resource condition of various devices. However, previous adaptive depth networks do not provide general principles and a formal explanation on why and which layers can be skipped, and, hence, their approaches are hard to be generalized and require long and complex training steps. In this paper, we present a practical approach to adaptive depth networks that is applicable to various networks with minimal training effort. In our approach, every hierarchical residual stage is divided into two sub-paths, and they are trained to acquire different properties through a simple self-distillation strategy. While the first sub-path is essential for hierarchical feature learning, the second one is trained to refine the learned features and minimize performance degradation if it is skipped. Unlike prior adaptive networks, our approach does not train every target sub-network in an iterative manner. At test time, however, we can connect these sub-paths in a combinatorial manner to select sub-networks of various accuracy-efficiency trade-offs from a single network. We provide a formal rationale for why the proposed training method can reduce overall prediction errors while minimizing the impact of skipping sub-paths. We demonstrate the generality and effectiveness of our approach with convolutional neural networks and transformers.
@@ 1 Introduction
There have been some attempts to provide predictable adaptability to neural networks by exploiting the redundancy in either network depths [8, 9], widths [10, 11], or both [12, 13]. However, one major difficulty with prior adaptive networks is that they are hard to train and require significantly longer training time than non-adaptive networks. For example, most adaptive networks select a fixed number of sub-networks of varying depths or width, and train them iteratively, mostly through self-distilling knowledge from the largest sub-network (also referred to as the super-net) [10, 11, 13, 14]. However, this exhaustive self-distillation takes long time and can generate conflicting training objectives for different parameter-sharing sub-networks, potentially resulting in worse performance [15, 16]. Moreover, unlike width-adaptation networks, there is no established principle for adapting depths since the impact of skipping individual layers has not been formally defined.
@@
Figure 1: (a) During training, each residual stage of a network is divided into two sub-paths. The layers in the second sub-paths are optimized to minimize performance degradation even if they are skipped. (b) At test time, we can instantly select $2^{N_{r}}$ parameter-sharing sub-networks from a single network by systematically skipping the last half layers of $N_{r}$ residual stages. (c) The sub-networks selected from a single network form a better Pareto frontier compared to individual networks.
@@
In this work, we introduce an architectural pattern and training method for adaptive depth networks that is generally applicable to various networks, e.g., CNNs and transformers. In the proposed adaptive depth networks, every residual stage is divided into two sub-paths and the sub-paths are trained to have different properties. While the first sub-paths are mandatory for hierarchical feature learning, the second sub-paths are optimized to incur minimal performance degradation even if they are skipped. In order to enforce this property of the second sub-paths, we propose a simple self-distillation strategy, in which only the largest sub-network (or, super-net) and the smallest sub-network (or, base-net) are exclusive used as a teacher and a student, respectively, as shown in Figure-1-(a). The proposed self-distillation strategy does not require exhaustive training of every target sub-network, resulting in significantly shorter training time than prior adaptive networks. Nevertheless, at test time, sub-networks with various depths can be selected instantly from a single network by connecting these sub-paths in a combinatorial manner, as shown in Figure 1-(b). Further, these sub-networks with varying depths outperform individually trained non-adaptive networks due to the regularization effect, as shown in Figure 1-(c).
@@
In Section 3, we discuss the details of our architectural pattern and training algorithm, and show formally that the selected sub-paths trained with our self-distillation strategy are optimized to reduce prediction errors while minimally changing the level of input features. In Section 4, we empirically demonstrate that our adaptive depth networks outperform counterpart individual networks, both in CNNs and vision transformers, and achieve actual inference acceleration and energy-saving.
To authors’ best knowledge, this is the first approach to adaptive networks that provides a general principle and theoretic basis for how depth adaptation can be supported with minimal performance degradation and training effort. This allows us to avoid typical exhaustive training of target sub-networks and instead construct sub-networks of varying depths from specifically trained sub-paths. We anticipate these advances will facilitate future research and applications of adaptive networks.
@@ 2 Related Work
Adaptive Networks: In most adaptive networks, parameter-sharing sub-networks are instantly selected by adjusting either widths, depths, or resolutions [8, 10, 11, 17, 13, 15, 18, 12, 19]. For example, slimmable neural networks adjust channel widths of CNN models on the fly for accuracy-efficiency trade-offs and they exploit switchable batch normalization to handle multiple sub-networks [10, 20, 11]. Transformer-based adaptive depth networks have been proposed for language models to dynamically skip some of the layers during inference [13, 9]. However, in these adaptive networks, every target sub-network with varying widths or depths need to be trained explicitly, incurring significant training overheads and potential conflicts between sub-networks.
Dynamic Networks: Dynamic networks [21] are another class of adaptive networks that exploit additional control networks or decision gates for input-dependent adaptation of CNN models [22, 23, 24, 25, 26] and transformers [27, 28, 29, 30]. In particular, most dynamic networks for depth-adaptation have some kinds of decision gates at every layers (or blocks) that determine if the layers can be skipped [31, 22, 32, 33]. These approaches are based on the thought that some layers can be skipped on ‘easy’ inputs. However, the learned policy for skipping layers is opaque to users and does not provide a formal description of which layers can be skipped for a given input. Therefore, the network depths cannot be adapted in a predictable manner to meet the resource condition of target devices.
Residual Blocks with Shortcuts: Since the introduction of ResNets [34], residual blocks with identity shortcuts have received extensive attention because of their ability to train very deep networks, and have been chosen by many CNNs [35, 36] and transformers [1, 37, 38]. Veit et al. [39] argue that identity shortcuts make exponential paths and results in an ensemble of shallower sub-networks. This thought is supported by the fact that removing individual residual blocks at test time does not significantly affect performance, and it has been further exploited to train deep networks [40, 41]. Other works argue that identity shortcuts enable residual blocks to perform iterative feature refinement, where each block improves slightly but keeps the semantic of the representation of the previous layer [42, 43]. Our work build upon those views on residual blocks with shortcuts and further extend them for adaptive depth networks by introducing a training method that enforces the properties of residual blocks more explicitly for skippable sub-paths.
@@ 3 Adaptive Depth Networks
We first present the architecture and training details of adaptive depth networks. Then, we discuss the theoretic rationale for how depth adaptation can be achieved with minimal performance degradation.
Figure 2: Illustration of a residual stage with two sub-paths. While the first (blue) sub-path is mandatory for hierarchical feature learning, the second (orange) sub-path can be skipped for efficiency. The layers in the skippable sub-path are trained to preserve the feature distribution from $\mathbf{h}^{s}_{base}$ to $\mathbf{h}^{s}_{super}$ using the proposed self-distillation strategy. Having similar feature distributions, either $\mathbf{h}^{s}_{base}$ or $\mathbf{h}^{s}_{super}$ can be provided as an input $\mathbf{h}^{s}$ to the next residual stage. In the mandatory sub-path, another set of batch normalization operators, called skip-aware BNs, is exploited when the second sub-path is skipped. This architectural pattern is applicable to both CNNs and transformers.
@@ 3.1 Architectural Pattern for Depth Adaptation
In hierarchical networks, such as ResNets [34] and Swin transformers [38], there are typically 4 to 7 consecutive residual stages. In a network with $N_{r}$ residual stages, the $s$-th ($s=1,...,N_{r}$) residual stage consists of $L$ identical residual blocks  [^1] In transformers, residual blocks are called encoder-decoder blocks. In most networks, $L$ is usually 3 to 6. that transform input features $\mathbf{h}^{s}_{1}$ additively to produce the output features $\mathbf{h}^{s}$, as follows:
$\underbrace{\mathbf{h}^{s}_{1}}_{=\mathbf{h}^{s-1}}+\underbrace{F_{1}(\mathbf{h}^{s}_{1})+...+{F}^{s}_{L/2}(\mathbf{h}^{s}_{L/2})}_{\mathbf{F}^{s}_{base}}+\underbrace{F_{L/2+1}(\mathbf{h}^{s}_{L/2+1})...+F_{L}(\mathbf{h}^{s}_{L})}_{\mathbf{F}^{s}_{skippable}}=\mathbf{h}^{s}$ (1)
@@
While a block with a residual function $F_{\ell}$ ($\ell=1,...,L$) learns hierarchical features as traditional compositional networks [44], previous literature [43, 42] demonstrates that a residual function also tend to learn a function that refines already learned features at the same feature level. If a residual block mostly performs feature refinement while not changing the level of input features, the performance of the residual network is not significantly affected by dropping the block at test time [40, 41]. However, in typical residual networks, most residual blocks tend to refine features while learning new level features as well, and, hence, random dropping of residual blocks at test time degrades inference performance significantly. Therefore, we hypothesize that if some selected residual blocks can be encouraged explicitly during training to focus more on feature refinement, then these blocks can be skipped to save computation at marginal loss of prediction accuracy at test time.
@@
To this end, we propose an architectural pattern for adaptive depth networks, in which every residual stage is divided into two sub-paths, or $\mathbf{F}^{s}_{base}$ and $\mathbf{F}^{s}_{skippable}$ as in Equation 1 and Figure 2. We train these two sub-paths to have different properties. While $\mathbf{F}^{s}_{base}$ is trained to learn feature representation $\mathbf{h}^{s}_{base}$ (= $\mathbf{h}^{s}_{L/2+1}$) with no constraint, the second sub-path $\mathbf{F}^{s}_{skippable}$ is constrained to preserve the feature level of $\mathbf{h}^{s}_{base}$ and only refine it to produce $\mathbf{h}^{s}_{super}(=\mathbf{h}^{s}_{L+1}$). Since layers in $\mathbf{F}^{s}_{base}$ perform essential transformations for hierarchical feature learning, they cannot be bypassed during inference. In contrast, layers in $\mathbf{F}^{s}_{skippable}$ can be skipped to save computation since they only refine $\mathbf{h}^{s}_{base}$. If $\mathbf{F}^{s}_{skippable}$ is skipped, then intermediate features $\mathbf{h}^{s}_{base}$ becomes the input to the next residual stage. Therefore, $2^{N_{r}}$ sub-networks with different accuracy-efficiency trade-offs can be selected from a single network by choosing whether to skip $\mathbf{F}^{s}_{skippable}(s=1,...,N_{r})$ or not.
Table 4 in Appendix A.1 provides details on how this pattern is applied to both CNNs and transformers.
3.2 Training Sub-Paths with Self-Distillation
@@
Preserving the feature level of $\mathbf{h}^{s}_{base}$ in $\mathbf{F}^{s}_{skippable}$ implies, more specifically, that two feature representations $\mathbf{h}^{s}_{base}$ and $\mathbf{h}^{s}_{super}$ have similar distributions over training input $\mathbf{X}$. Algorithm 1 shows our training method, in which $\mathbf{h}^{s}_{base}$ and $\mathbf{h}^{s}_{super}$ are encouraged to have similar distributions by including Kullback-Leibler (KL) divergence between them, or ${D_{KL}}(\mathbf{h}^{s}_{super}||\mathbf{h}^{s}_{base})$, in the loss function.
Algorithm 1 Training algorithm for an adaptive depth network $\mathbf{M}$. The forward function of $\mathbf{M}$ accepts an extra argument, ‘skip’, which controls the residual stages where their skippable sub-paths are skipped. For example, the smallest sub-network, or base-net, of $\mathbf{M}$ is selected by passing ‘skip=[True, True, True, True]’ when the total number of residual stages, denoted by $N_{r}$, is 4.
@@
9:     $loss_{base}=\sum_{s=1}^{N_{r}}D_{KL}(\mathbf{h}^{s}_{super}\|\mathbf{h}^{s}_{base})+D_{KL}(\mathbf{\hat{y}}_{super}\|\mathbf{\hat{y}}_{base})$
10:     $loss_{base}.backward()$triangleright self-distillation of skippable sub-paths
@@
In Algorithm 1, the largest and the smallest sub-networks of $\mathbf{M}$, which are called super-net and the base-net, respectively, are exploited. During the forward passes of the super-net and the base-net in steps 5 and 8, respectively, intermediate features $\mathbf{h}_{super}$ and $\mathbf{h}_{base}$ are obtained. In steps 9-10, $D_{KL}(\mathbf{h}^{s}_{super}||\mathbf{h}^{s}_{base})$ in the loss function is explicitly minimized for training data $\mathbf{X}$. Further, this self-distillation step gives the effect of transferring the knowledge from $\mathbf{h}^{s}_{super}$ to $\mathbf{h}^{s}_{base}$ at every residual stage. Therefore, $\mathbf{h}^{s}_{base}$ is expected to learn more compact representation from $\mathbf{h}^{s}_{super}$.
Self-distillation itself has been extensively used in prior adaptive networks [11, 10, 13]. However, their primary goal is to train every target sub-network. For example, on each iteration, sub-networks are randomly sampled from a large search space to act as either teachers or students [11], which results in significantly longer training time. In contrast, our method in Algorithm 1 focuses on training sub-paths by exclusively using the super-net as a teacher and the base-net as a student. By focusing on training sub-paths rather than target sub-networks, the training procedure in Algorithm 1 is significantly simplified. At test time, however, $2^{N_{r}}(=\sum_{k=0}^{N_{r}}{C(N_{r},k)})$ different sub-networks can be constructed instantly by connecting these sub-paths in a combinatorial manner. The effect of this self-distillation strategy is investigated in Section 4.4.
3.3 Analysis of Skippable Sub-Paths
Formal Analysis: $D_{KL}(\mathbf{h}^{s}_{super}||\mathbf{h}^{s}_{base})$ in the loss function $loss_{base}$ can be trivially minimized if residual blocks in $\mathbf{F}^{s}_{skippable}$ learn identity functions, or $\mathbf{h}^{s}_{base}+\mathbf{F}^{s}_{skippable}(\mathbf{h}^{s}_{base})=\mathbf{h}^{s}_{base}$. However, since the super-net is jointly trained with the loss function $loss_{super}$, the residual functions in $\mathbf{F}^{s}_{skippable}$ cannot simply be an identity function. Then, what do the residual functions in $\mathbf{F}^{s}_{skippable}$ learn during training? This can be further investigated through Taylor expansion [43]. For our adaptive depth networks, a loss function $\mathcal{L}$ used for training the super-net can be approximated with Taylor expansion as follows: